OpenStudy (summersnow8):

Solving for Km: I understand how I find Vmax, but I am having trouble solving for Kmax.....

OpenStudy (summersnow8):

I was thinking that slope = Km / Vmax Km = slope x Vmax Km = 7 x 10 ^-8 * (1/0.0071) Km = 9.86 x 10^-6

OpenStudy (summersnow8):

but I know that - 1 / Km so is Km actually -1/9.86 x 10^-6 = - 1.0 x 10^-5

OpenStudy (frostbite):

Hey @Summersnow8 You are correct in your calculation of \(K_{M}\) through the use of the slope in your Lineweaver-Burk plot. The \(\frac{-1}{K_M}\) is equal to the value of the intercept with the x-axis and should equal the same.

OpenStudy (frostbite):

I have done so much enzyme kinetics so if anything please just ask ahead. :)